OpenStudy (anonymous):

Which is the value of n when 6n - 10 = 4n + 6 ?

OpenStudy (anonymous):

Show the work and the answer please

OpenStudy (haseeb96):

for isolation of n we will bring 4n on the left hand side and 10 on the righthand side and change their sign so the equation becomes 6n -4n = 6 +10 2n =16 n=16/2 n=8 is the correct answer

OpenStudy (anonymous):

Ty

OpenStudy (haseeb96):

@MikaylaSkyy if u want to check either my value of n is correct or incorrect just put it in the above equation in the place of n and when u will find that your both hand side value is same then answer is correct and when not same tthen answer is incorrect

OpenStudy (haseeb96):

6n-10 = 4n + 6 now put n=8 6(8)-10 =4(8)+6 48 -10 = 32 +6 38 =38 so it is showing that answer is correct